# Floquet Nonadiabatic Mixed Quantum-Classical Dynamics in Laser-Dressed Solid Systems

### v1

@@ Abstract
In this paper, we introduce the Floquet Ehrenfest and Floquet surface hopping approaches to study the nonadiabatic dynamics in the laser-dressed solid systems. We demonstrate that these two approaches can be formulated in both real and reciprocal spaces. Using these approaches, we are able to simulate the interaction between electronic carriers and phonons under periodic drivings, such as strong light-matter interactions. Employing the Holstein and Peierls models, we show that the strong light-matter interactions can effectively modulate the dynamics of electronic population and mobility. Notably, our study demonstrates the feasibility and effectiveness of modeling low-momentum carriers’ interactions with phonons using a truncated reciprocal-space basis, an approach impractical in real-space frameworks. Moreover, we reveal that even with significant truncation, carrier populations derived from surface hopping maintain greater accuracy compared to those obtained via meanfield dynamics. These results underscore the potential of our proposed methods in advancing the understanding of carrier-phonon interactions in various laser-dressed materials.
@@ I Introduction
Nowadays, strong light-matter interaction have become prominent, leading to significant and observable effects. With the development of modern techniques, properties of solids have been modified by light-matter strong coupling or laser fields Garcia-Vidal, Ciuti, and Ebbesen (2021). For example, in the field of quantum computation, researchers have proposed a scheme to accelerate the nontrivial two-qubit phase gate by ultrastrongly coupling superconducting flux qubits to a circuit quantum electrodynamics systemWang et al. (2017). Experiments have demonstrated that materials subjected to intense terahertz pulses exhibit transient superconducting properties at significantly higher temperatures than at equilibrium Mankowsky et al. (2014); Yang et al. (2019). Applying electromagnetic radiation to a topologically trivial insulator, such as a non-inverted HgTe/CdTe quantum well with no edge state in the static limit, can induce novel topological edge states Lindner, Refael, and Galitski (2011). Suitably chosen irradiation parameters can open gaps at the Dirac point of graphene and even drive graphene into the topological Haldane phase Kitagawa et al. (2010, 2011).
@@
From theoretical perspective, the analysis of periodically driven systems often relies on Floquet theorem Bukov, D’Alessio, and Polkovnikov (2015). Floquet formalism allows for the reduction of the periodic or quasi-periodical time-dependent Hamiltonian into time-independent Floquet matrix Chu and Telnov (2004). Several theoretical methods have been developed under the Floquet representation to solve problems related to periodically driven solid states. For example, Kitagawa $et$ $al$. Kitagawa et al. (2011) demonstrated that quantum transport properties can be controlled in materials such as graphene and topological insulators by applying light using Floquet Green’s function. Tsuji $et$ $al$. Tsuji, Oka, and Aoki (2008) studied correlated electron systems that are periodically driven by external fields via Floquet dynamical mean-field theory and showed that photoinduced midgap states emerge from strong ac fields, triggering an insulator-metal transition.
@@
In this study, we develop two trajectory-based nonadiabatic mix quantum-classical dynamics methods, namely the Floquet meanfield dynamics method (FMD) and the Floquet surface hopping (FSH) algorithm, for solid states under external periodic drivings. We consider interactions between electronic carriers and phonons both in real and reciprocal spaces. Real space simulations are generally more suitable for disordered solids, while reciprocal space simulations are more suitable for ordered solids. Previously, several theoretical methods have been proposed to describe lattice-based electron-phonon couplings beyond perturbative regimes, such as mixed quantum-classical dynamicsKrotz, Provazza, and Tempelaar (2021), and surface hoppingKrotz and Tempelaar (2022); Xie et al. (2022). Using a tight-binding model involving Holstein-type Holstein (1959); Freericks, Jarrell, and Scalapino (1993); Zhang, Jeckelmann, and White (1999) and Peierls-type Su, Schrieffer, and Heeger (1979) electron-phonon couplings, we calculate electronic dynamics and charge mobility under different drivings. We expect that the FMD and FSH algorithms, both in real- and reciprocal-space, can provide theoretical predictions on solid state property modified by external periodic drivings.
@@ II.1 Model Hamiltonian
For a one-dimensional lattice of harmonic, noninteracting, and dispersionless phonons, the real space Hamiltonian is given by:
@@
where $\omega$ indicates the phonon energy. The transformation from real to reciprocal space gives as:
@@
we can get the position $\tilde{r}_{k}$ and momentum $\tilde{p}_{k}$ in the reciprocal space as:
@@
Note that under the transformation from real to reciprocal space, classical position and momentum coordinates become scrambled. Then, we obtain the phonon Hamiltonian in the reciprocal space:
@@
where $\hat{c}_{n}^{(\dagger)}$ and $\hat{c}_{k}^{(\dagger)}$ are real- and reciprocal-space electronic annihilation (creation) operators, respectively. $(J_{0}+J_{1}\cos(\Omega t))$ is the nearest-interaction term. Here, we assume that lasers can affect the hopping probability between two sides. $J_{0}$ signifies the coupling strength inherent to the electronic states, while $J_{1}\cos(\Omega t)$ provides a semi-classical portrayal of light-matter interactions: $J_{1}$ is the coupling strength, and $\Omega$ is the driving frequency. By placing this term in off-diagonal terms, we highlight how light influences the couplings among different electron states.
Then we give the Hamiltonian of electron-phonon coupling $\hat{H}_{\text{el-ph}}$ for two models. Firstly, as for the Holstein model:
@@ II.2 Floquet Theory
Then we can expand our wavefunction in the following form, which is commonly used to describe energy eigenstates and their evolution:
@@
$&u^{\dagger}(\mathbf{R},t)H^{\text{F}}(\mathbf{R},t)u(\mathbf{R},t)=\sum_{\beta=-m_{\text{F}}}^{m_{\text{F}}}\sum_{\alpha=-n_{\text{F}}}^{n_{\text{F}}}\\ &(c^{\dagger}_{\beta}(r)e^{-i\beta\Omega t}\hat{H}(\mathbf{R},t)e^{i\alpha\Omega t}c_{\alpha}(r)-c^{\dagger}_{\beta}(r)c_{\alpha}(r)\alpha\hbar\Omega e^{i(\alpha-\beta)\Omega t})\end{aligned}$ (15)
@@
$\begin{split}[(\sum_{\beta=-m_{\text{F}}}^{m_{\text{F}}}\sum_{\alpha=-n_{\text{F}}}^{n_{\text{F}}}&\frac{1}{T}\int_{0}^{T}dte^{-i\beta\Omega t}\hat{H}(\mathbf{R},t)e^{i\alpha\Omega t})\\ &-\alpha\hbar\Omega\delta_{\alpha\beta}]c_{\alpha}(r)=\epsilon c_{\beta}(r).\end{split}$ (16)
Finally, we can get the time-independent Floquet Hamiltonian:
$\hat{H}^{\text{F}}(\mathbf{R})=\sum_{\alpha}H_{\beta\alpha}(\mathbf{R})-\alpha\hbar\Omega\delta_{\alpha\beta},$ (17)
Now, we can transform the Hamiltonian into the Floquet representation. As for the Holstein model, we can construct each single block $H_{\beta\alpha}$ as:
$\begin{split}H_{\beta\alpha}=&[-J_{0}\sum_{n}\left(\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+\text{h.c.}\right)+g\sqrt{2\omega^{3}}\sum_{n}\hat{c}_{n}^{\dagger}\hat{c}_{n}r_{n}\delta_{\alpha\beta}\\ &-\frac{1}{2}J_{1}\sum_{n}\left(\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+\text{h.c.}\right)(\delta_{\alpha,\beta+1}+\delta_{\alpha,\beta-1})\\ =&[-2J_{0}\sum_{k}\hat{c}_{k}^{\dagger}\hat{c}_{k}\cos(k)+g\sqrt{\frac{\omega}{2N}}\sum_{k,\kappa}\hat{c}_{k+\kappa}^{\dagger}\hat{c}_{k}(\omega(\tilde{r}_{-\kappa}+\tilde{r}_{\kappa})\\ &-i(\tilde{p}_{-\kappa}-\tilde{p}_{\kappa}))]\delta_{\alpha\beta}-J_{1}\sum_{k}\hat{c}_{k}^{\dagger}\hat{c}_{k}\cos(k)(\delta_{\alpha,\beta+1}+\delta_{\alpha,\beta-1}),\end{split}$
@@
$H_{\beta\alpha}=&[-J_{0}\sum_{n}\left(\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+\mathrm{h.c.}\right)\\ &+g\sqrt{2\omega^{3}}\sum_{n}(\hat{c}_{n}^{\dagger}\hat{c}_{n+1}+\hat{c}_{n+1}^{\dagger}\hat{c}_{n})(r_{n}-r_{n+1})]\delta_{\alpha\beta}\\ &-\frac{1}{2}J_{1}\sum_{n}\left(\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+\text{h.c.}\right)(\delta_{\alpha,\beta+1}+\delta_{\alpha,\beta-1})\\ =&[-2J_{0}\sum_{k}\hat{c}_{k}^{\dagger}\hat{c}_{k}\cos(k)+g\sqrt{\frac{2\omega}{N}}\sum_{k,\kappa}\hat{c}_{k+\kappa}^{\dagger}\hat{c}_{k}(i\omega(\tilde{r}_{-\kappa}+\tilde{r}_{\kappa})\\ &+(\tilde{p}_{-\kappa}-\tilde{p}_{\kappa}))(\sin(k+\mathbf{\kappa})-\sin(k))]\delta_{\alpha\beta}\\ &-J_{1}\sum_{k}\hat{c}_{k}^{\dagger}\hat{c}_{k}\cos(k)(\delta_{\alpha,\beta+1}+\delta_{\alpha,\beta-1}).\end{aligned}$
@@
Figure 1: Floquet replicas without electron-phonon coupling in reciprocal space. Here, we set the Floquet levels $n_{\text{F}}=2$, so there are $5$ quasi-surfaces in total. Other parameters: $J=0.3$, $\hbar\omega=0.3$, $g=0.5$, and $kT=0.5$.
@@ II.3 Floquet Meanfield Dynamics (FMD)
In this subsection, we will introduce the Floquet meanfield approach in real and reciprocal space. Firstly, according to Schrödinger’s equation, we can propagate the wavefunction as:
@@
Here, $\hat{H}^{\text{F}}(\mathbf{R})$ including pure electron Hamiltonian and electron-phonon coupling Hamiltonian. We can express the Ehrenfest dynamics in real space as follows:
@@ II.4 Floquet Surface Hopping(FSH)
Within the surface hopping algorithm, one propagates a trajectory along the adiabatic surface, and integrates the electronic density matrix. Here, we propagate a trajectory along the Floquet adiabatic surface. The hopping rate from surface $J$ to surface $I$ can be determined bySubotnik et al. (2016); Wang, Akimov, and Prezhdo (2016)
@@
where $\Theta$ is the Heaviside function
$\Theta(x)=\begin{cases}x&\text{x>0}\\ 0&\text{x<0}\end{cases}$ (22)
@@
Here, we set nuclear mass and $\hbar$ equal to $1$. $\hat{\rho}^{\text{F(ad)}}=\ket{\Psi^{\text{F(ad)}}}\bra{\Psi^{\text{F(ad)}}}$ | is the density matrix in adiabatic representation. $\bm{d}_{IJ}$ is the derivative coupling
@@
After a hop from Floquet state $J$ to Floquet state $I$, one should rescale the nuclear momentum in the direction of derivative coupling $\bm{d}_{JI}$ubotnik, Ouyang, and Landry (2013); Dou and Subotnik (2017)
@@
where the $\xi$ can be solved by energy conservation
@@
where $\xi$ can be solved by energy conservation,

### v2

@@ Abstract
In this paper, we introduce the Floquet Ehrenfest and Floquet surface hopping approaches to study the nonadiabatic dynamics in the laser-dressed solid systems. We demonstrate that these two approaches can be formulated in both real and reciprocal spaces. Using the two approaches, we are able to simulate the interaction between electronic carriers and phonons under periodic drivings, such as strong light-matter interactions. Employing the Holstein and Peierls models, we show that the strong light-matter interactions can effectively modulate the dynamics of electronic population and mobility. Notably, our study demonstrates the feasibility and effectiveness of modeling low-momentum carriers’ interactions with phonons using a truncated reciprocal space basis, an approach impractical in real space frameworks. Moreover, we reveal that even with significant truncation, carrier populations derived from surface hopping maintain greater accuracy compared to those obtained via meanfield dynamics. These results underscore the potential of our proposed methods in advancing the understanding of carrier-phonon interactions in various laser-dressed materials.
@@ I Introduction
Nowadays, strong light-matter interactions have become prominent, leading to significant and observable effects. With the development of modern techniques, properties of solids have been modified by light-matter strong couplings or laser fields Garcia-Vidal, Ciuti, and Ebbesen (2021). For example, in the field of quantum computation, researchers have proposed a scheme to accelerate the nontrivial two-qubit phase gate by ultrastrongly coupling superconducting flux qubits to a circuit quantum electrodynamics systemWang et al. (2017); Schirò, Eich, and Agostini (2021); Cabra, Franco, and Galperin (2020); Tiwari, Gu, and Franco (2023). Experiments have demonstrated that materials subjected to intense terahertz pulses exhibit transient superconducting properties at significantly higher temperatures than at equilibrium Mankowsky et al. (2014); Yang et al. (2019). Applying electromagnetic radiation to a topologically trivial insulator, such as a non-inverted HgTe/CdTe quantum well with no edge state in the static limit, can induce novel topological edge states Lindner, Refael, and Galitski (2011). Suitably chosen irradiation parameters can open gaps at the Dirac point of graphene and even drive graphene into the topological Haldane phase Kitagawa et al. (2010, 2011).
@@
From theoretical perspective, the analysis of periodically driven systems often relies on Floquet theorem Bukov, D’Alessio, and Polkovnikov (2015). Floquet formalism allows for the reduction of the periodic or quasi-periodical time-dependent Hamiltonian into time-independent Floquet matrix Chu and Telnov (2004); Mori (2023); Sato et al. (2020). Several theoretical methods have been developed under the Floquet representation to solve problems related to periodically driven solid states. For example, Kitagawa $et$ $al$. Kitagawa et al. (2011) demonstrated that quantum transport properties can be controlled in materials such as graphene and topological insulators by applying light using Floquet Green’s function. Tsuji $et$ $al$. Tsuji, Oka, and Aoki (2008) studied correlated electron systems that are periodically driven by external fields via Floquet dynamical mean-field theory and showed that photoinduced midgap states emerge from strong ac fields, triggering an insulator-metal transition.
In this study, we develop two trajectory-based nonadiabatic mix quantum-classical dynamics methods, namely the Floquet meanfield dynamics method (FMD) and the Floquet surface hopping (FSH) algorithm, for solid states under external periodic drivings. Our research builds upon the foundation laid by existing real and reciprocal space semi-classical meanfield dynamics and surface hopping methods, as documented in Refs. [22, 23]. Furthermore, our approach to Floquet mixed quantum-classical dynamics is directly connected to the Floquet quantum-classical Liouville equation, which has been rigorously derived in Ref. [24]. Note also that the connection between quantum-classical Liouville equation and mix quantum-classical dynamics, such as surface hopping has been established in Ref. [25]. While Floquet surface hopping has been employed to investigate mixed quantum-classical dynamics under periodic driving in the gas phase, see Ref. [26]; our current study focuses on implementing FSH and FMD for solids.
@@
We consider the interactions between electronic carriers and phonons both in real and reciprocal spaces. Real space simulations are generally more suitable for disordered solids, while reciprocal space simulations are more suitable for ordered solids. Using a tight-binding model involving Holstein-type Holstein (1959); Freericks, Jarrell, and Scalapino (1993); Zhang, Jeckelmann, and White (1999) and Peierls-type Su, Schrieffer, and Heeger (1979) electron-phonon couplings, we calculate electronic dynamics and charge mobility under different drivings. We expect that the FMD and FSH algorithms, both in real and reciprocal space, can provide theoretical predictions on solid state property modified by external periodic drivings.
@@ II.1 Model Hamiltonian
For a one-dimensional lattice of harmonic, noninteracting, and dispersionless phonons Rosati, Dolcini, and Rossi (2015); Glebov et al. (2000), the real space Hamiltonian is given by:
@@
where $\omega$ indicates the phonon energy. We remove the hat for $H_{\text{ph}}$ since we make classical approximation here. Note that our methods developed here are not restricted to one phonon frequency. Our methods can be used to study multiple phonons with different frequencies easily. The transformation from real to reciprocal space gives as:
@@
we can get the position $\tilde{r}_{k}$ and momentum $\tilde{p}_{k}$ in the reciprocal space as Krotz, Provazza, and Tempelaar (2021); Krotz and Tempelaar (2022):
@@
Note that under the transformation from real to reciprocal space, classical position and momentum coordinates become scrambled. Then, we obtain the phonon Hamiltonian in the reciprocal space with classical approximation:
@@
where $\hat{c}_{n}^{(\dagger)}$ and $\hat{c}_{k}^{(\dagger)}$ are real and reciprocal space electronic annihilation (creation) operators, respectively. $(J_{0}+J_{1}\cos(\Omega t))$ is the nearest-interaction term. Note the light-matter interactions is introduced where the light couples to transition dipole between two sites: $J_{0}$ signifies the coupling strength inherent to the electronic states, while $J_{1}\cos(\Omega t)$ provides a semi-classical portrayal of light-matter interactions: $J_{1}$ is the coupling strength, and $\Omega$ is the driving frequency. This coupling has been investigated in Refs. [33; 34], which can be achieved in experiments, for example, by periodically bending an ultrathin metallic array of coupled corrugated waveguides Cheng et al. (2019).
Then we give the Hamiltonian of electron-phonon coupling $\hat{H}_{\text{el-ph}}$ for two models. Firstly, as for the Holstein model Krotz, Provazza, and Tempelaar (2021); Krotz and Tempelaar (2022):
@@ II.2 Floquet Theory
Then we can expand our wavefunction in the following form as one solution to Eq. (11), which is commonly used to describe energy eigenstates and their evolution:
@@
$&u^{\dagger}(\mathbf{R},t)H^{\text{F}}(\mathbf{R},t)u(\mathbf{R},t)=\sum_{\beta=-m_{\text{F}}}^{m_{\text{F}}}\sum_{\alpha=-n_{\text{F}}}^{n_{\text{F}}}\\ &(c^{\dagger}_{\beta}(\mathbf{R})e^{-i\beta\Omega t}\hat{H}(\mathbf{R},t)e^{i\alpha\Omega t}c_{\alpha}(\mathbf{R})-c^{\dagger}_{\beta}(\mathbf{R})c_{\alpha}(\mathbf{R})\alpha\hbar\Omega e^{i(\alpha-\beta)\Omega t}).\end{aligned}$ (15)
@@
$\begin{split}[(\sum_{\beta=-m_{\text{F}}}^{m_{\text{F}}}\sum_{\alpha=-n_{\text{F}}}^{n_{\text{F}}}&\frac{1}{T}\int_{0}^{T}dte^{-i\beta\Omega t}\hat{H}(\mathbf{R},t)e^{i\alpha\Omega t})\\ &-\alpha\hbar\Omega\delta_{\alpha\beta}]c_{\alpha}(\mathbf{R})=\epsilon c_{\beta}(\mathbf{R}).\end{split}$ (16)
The procedure described in Eq. (16) corresponds to representing the equation of motion in Floquet space, where we can define the Floquet Hamiltonian:
$\hat{H}^{\text{F}}(\mathbf{R})=\sum_{\alpha}H_{\beta\alpha}(\mathbf{R})-\alpha\hbar\Omega\delta_{\alpha\beta}.$ (17)
Note that the Hamiltonian in Floquet representation is time-independent. The time-dependency is incorporated in the replicas in the Floquet space. Note also that one can obtain the averaged observable directly in the Floquet representation (e.g., see Eqs. (21,22)).
Now, we can transform our model Hamiltonian into the Floquet representation. The Floquet theory used in this paper is a classical theory of light, which has the connection with optical cavity and can be used to simulate quantum electrodynamics Hübener et al. (2021). As for the Holstein model, we can construct each single block $H_{\beta\alpha}$ as:
$H_{\beta\alpha}=$ $[-J_{0}\sum_{n}\left(\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+\text{h.c.}\right)+g\sqrt{2\omega^{3}}\sum_{n}\hat{c}_{n}^{\dagger}\hat{c}_{n}r_{n}\delta_{\alpha\beta}$ (18)
$-\frac{1}{2}J_{1}\sum_{n}\left(\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+\text{h.c.}\right)(\delta_{\alpha,\beta+1}+\delta_{\alpha,\beta-1})$
$=$ $[-2J_{0}\sum_{k}\hat{c}_{k}^{\dagger}\hat{c}_{k}\cos(k)+g\sqrt{\frac{\omega}{2N}}\sum_{k,\kappa}\hat{c}_{k+\kappa}^{\dagger}\hat{c}_{k}(\omega(\tilde{r}_{-\kappa}+\tilde{r}_{\kappa})$
$-i(\tilde{p}_{-\kappa}-\tilde{p}_{\kappa}))]\delta_{\alpha\beta}-J_{1}\sum_{k}\hat{c}_{k}^{\dagger}\hat{c}_{k}\cos(k)(\delta_{\alpha,\beta+1}+\delta_{\alpha,\beta-1}),$
@@
$H_{\beta\alpha}=$ $[-J_{0}\sum_{n}\left(\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+\mathrm{h.c.}\right)$ (19)
$+g\sqrt{2\omega^{3}}\sum_{n}(\hat{c}_{n}^{\dagger}\hat{c}_{n+1}+\hat{c}_{n+1}^{\dagger}\hat{c}_{n})(r_{n}-r_{n+1})]\delta_{\alpha\beta}$
$-\frac{1}{2}J_{1}\sum_{n}\left(\hat{c}_{n+1}^{\dagger}\hat{c}_{n}+\text{h.c.}\right)(\delta_{\alpha,\beta+1}+\delta_{\alpha,\beta-1})$
$=$ $[-2J_{0}\sum_{k}\hat{c}_{k}^{\dagger}\hat{c}_{k}\cos(k)+g\sqrt{\frac{2\omega}{N}}\sum_{k,\kappa}\hat{c}_{k+\kappa}^{\dagger}\hat{c}_{k}(i\omega(\tilde{r}_{-\kappa}+\tilde{r}_{\kappa})$
$+(\tilde{p}_{-\kappa}-\tilde{p}_{\kappa}))(\sin(k+\mathbf{\kappa})-\sin(k))]\delta_{\alpha\beta}$
$-J_{1}\sum_{k}\hat{c}_{k}^{\dagger}\hat{c}_{k}\cos(k)(\delta_{\alpha,\beta+1}+\delta_{\alpha,\beta-1}).$
@@
Figure 1: Floquet replicas without electron-phonon coupling in reciprocal space. Here, we set the Floquet levels $n_{\text{F}}=2$, so there are $5$ quasi-surfaces in total. Other parameters: $J=0.3$, $\hbar\omega=0.3$, $g=0.5$, and $kT=0.5$. Here, the energy unit corresponds to $50.5$ $meV$.
@@ II.3 Floquet Meanfield Dynamics (FMD)
In this subsection, we will introduce the Floquet meanfield approach in real and reciprocal space. In Ref. [37], it has been demonstrated that the equation of motion of the density operator obeys Liouville-von Neumann equation in Floquet representation. Therefore, we can deduce the following equation:
@@
Here, $\hat{H}^{\text{F}}(\mathbf{R})$ including pure electron Hamiltonian and electron-phonon coupling Hamiltonian, while $\Psi^{\text{F}}$ represents the Floquet wavefunction. In Floquet space, each original surface has multiple Floquet replicas with energy gap of $\alpha\hbar\Omega$, where $\alpha$ is an integer ranging from negative infinity to positive infinity. Initially, we sample all trajectories on $\alpha=0$ Floquet state. We can express the Ehrenfest dynamics in real space as follows Krotz, Provazza, and Tempelaar (2021):
@@ II.4 Floquet Surface Hopping(FSH)
Within the surface hopping algorithm, one propagates a trajectory along the adiabatic surface, and integrates the electronic density matrix. Our FSH method is performed among multiple Floquet electronic orbitals, the hopping processes are elaborated in Ref. [23].Here, we propagate a trajectory along the Floquet adiabatic surface. The hopping rate from surface $J$ to surface $I$ can be determined bySubotnik et al. (2016); Wang, Akimov, and Prezhdo (2016); Krotz and Tempelaar (2022):
@@
where $\Theta$ is the Heaviside function:
$\Theta(x)=\begin{cases}x&\text{x>0}\\ 0&\text{x<0},\end{cases}$ (24)
@@
Here, we set nuclear mass and $\hbar$ equal to $1$. $\hat{\rho}^{\text{F(ad)}}=\ket{\Psi^{\text{F(ad)}}}\bra{\Psi^{\text{F(ad)}}}$ | is the density matrix in adiabatic representation. $\bm{d}_{IJ}$ is the derivative coupling:
@@
After a hop from Floquet state $J$ to Floquet state $I$, one should rescale the nuclear momentum in the direction of derivative coupling $\bm{d}_{JI}$ubotnik, Ouyang, and Landry (2013); Dou and Subotnik (2017):
@@
where the $\xi$ can be solved by energy conservation:
@@
where $\xi$ can be solved by energy conservation: